OpenStudy (fedorable):

Use a transformation to solve the equation. h + (–3) = –0.6 A. h = –3.6 B. h = –2.4 C. h = 2.4 D. h = 3.6

OpenStudy (anonymous):

http://www.geteasysolution.com/

OpenStudy (vheah):

Since you have a -3 and you want to isolate h by itself, you have to alter the sign and add 3 to both sides: h= 3+(-0.6) a positive sign * a negative sign = negative h= 3-0.6 h=2.4